OpenStudy (anonymous):

How do you solve this equation over the complex numbers? z^4+36=0 [SOLUTIN IN ALGEBRIC FORM]

OpenStudy (solomonzelman):

subtract 36 from both sides, and you get \[z^4=-36\]take the square root of both sides. \[z^2=\sqrt{36}~~~~~->~~~~~z^2=\sqrt{6} \times \sqrt{6} \times i~~~~~->~~~~~z^2=6i \] take the square root of both sides again. \[z=\sqrt{ 6i }\]

OpenStudy (solomonzelman):

Do you know what "i" is ?

OpenStudy (anonymous):

Yes, I do know. There are 4 different roots.. I have managed to get z^4=6i. Don't know what to do next.. need to use De Moiver?

OpenStudy (solomonzelman):

I forgot + -6i, and to then square both sides, sorry.

OpenStudy (solomonzelman):

I don't know what De movier is, sorry again

OpenStudy (anonymous):

\[ z= r e^{i t}\\ r^4 e^{4 i t} = 36 e^{ i \pi}\\\ r=\sqrt{6}\\ 4 t = \pi + 2 k \pi\\ t= \frac {\pi}4+ k \frac{\pi}2\\ t=\frac \pi 4\\ t=\frac \pi 4 + \frac \pi 2=3 \frac \pi 4\\ t=\frac \pi 4 + 2 \frac \pi 2=5 \frac \pi 4\\ t=\frac \pi 4 + 3 \frac \pi 2=7 \frac \pi 4\\ \] You replace each t and r by their values and you obtain the three roots

OpenStudy (anonymous):

four roots